equations of lines which passes through the points of intersection of the lines 4x-3y-1=0and 2x-5y+3=0and are equally inclined to the axes are:
y(+or-)x=0
y-1=(+or-)1(x-1)
x-1=(+or-)2(y-1)
none of these

Grade:11

1 Answers

Ajay
7 years ago
The Point of intersection is (1, 1)
The equation of line passing through this point is
y-1 = m(x-1) where m is slope
or y-mx + m-1 = 0.
Since line makes equal with both axes this is possible line makes 45 or 135 degree with x axis.
The slopes thus can be tan45 or tan 135 or m can take values 1  or -1
Putting m = 1 and -1 in equation we wll get two equations as follows
y = x and
y+x-2=0 as equations of lines